You are considering starting a fishing charter business in San Juan. The boat would cost $235,000 today and be depreciated straight line over 10 years. The purchase will be partially financed with a 10 year loan. At the end of the 8th year, the boat is to be sold at a salvage value and loan paid off (making scheduled payment in year-8 and also paying off remaining balance). Compute the IRR, NPV and Payback of the business FCFE using a straight line depreciation. Assume the full cost of the boat is depreciated over 10 years, so you must account for the net book value in year 8 when computing after-tax salvage.

Boat $235,000 T=0
Revenues $100,000 T=1-8
Operating Expenses $35,000 T=1-8 (Gas, Bait, Maint, Insurance, etc.)
Tax Rate 35.00%
Depr. Life (straight line) 10
Boat Salvage $50,000 T=8
Discount Rate 20.00%
Loan Principal $100,000 Amortizing Loan
Loan Maturity (yrs)                       10.00
Loan APR 9%
Loan Payment Freq Annual Assume annual for simplicity (otherwise would need to create monthly FCF forecast)
NWC % Revenues 5% Immediate, then liquidited by end of year 8 (T=8, NWC=0)
NPV ?
IRR ?
Payback ?
Answer #1

For calculating FCFE, we first need to calculate the Depreciation, Loan Schedule, Working Capital and Profit and Loss Account. These are calculated below:

Depreciation Schedule

Depreciation per Year = $ 235000 / 10 years = $ 23,500 per year ....striaght line

Net Book Value = Net Book Value (last year) - Depreciation

Year ending 0 1 2 3 4 5 6 7 8
Depreciation (straight line) 23500 23500 23500 23500 23500 23500 23500 23500
Net Book Value of Boat 211500 188000 164500 141000 117500 94000 70500 47000

Loan Schedule

Repayment: Equal amortization assumed i.e. $ 10,000 repaid every year.

Loan Closing = Opening Loan - Repayment

Interest = 9% * Opening Loan

Year ending 0 1 2 3 4 5 6 7 8
Opening Loan 100000 90000 80000 70000 60000 50000 40000 30000
Less: Repayment 10000 10000 10000 10000 10000 10000 10000 30000
Loan Closing 100000 90000 80000 70000 60000 50000 40000 30000 0
Interest @9% APR 9000 8100 7200 6300 5400 4500 3600 2700

Working Capital

Working Capital = 5% of Revenues = 5% * $1000,000

Working Capital = $ 5,000

The revenues are stable, hence, there will be no change in working capital.

In year 8, when the operation closes, the working capital shall be released.

Year ending 0 1 2 3 4 5 6 7 8
Net Working Capital 5000 5000 5000 5000 5000 5000 5000 0
Change in Working Capital 5000 0 0 0 0 0 0 -5000

Salvage Value

Net Book value at end of Year 8 = $ 47,000 ....as per depreciation schedule

Salvage Value = $ 50,000

Hence, at the end of year 8, a profit of $ 3,000 (50,000 - 47,000) will be booked. The net salvage value is $3,000.

Based on the above Depreciation, Loan Schedule, Net Salvage Value and Working Capital Cycle, the Profit and Loss Account is given below:

P&L Account

Year Ending 0 1 2 3 4 5 6 7 8
Revenues 100000 100000 100000 100000 100000 100000 100000 100000
Less: Operating Expenses 35000 35000 35000 35000 35000 35000 35000 35000
EBITDA 65000 65000 65000 65000 65000 65000 65000 65000
Less: Depreciation 23500 23500 23500 23500 23500 23500 23500 23500
Less: Interest 9000 8100 7200 6300 5400 4500 3600 2700
Add: Net Salvage Value 3000
Profit Before Tax 32500 33400 34300 35200 36100 37000 37900 41800
Less: Taxes 11375 11690 12005 12320 12635 12950 13265 14630
Net Income 21125 21710 22295 22880 23465 24050 24635 27170

Now, the FCFE Can be calculated as:

FCFE = Net Income + Non Cash Expenses (Depreciation) - Change in Working Capital - Capital Expenditure + Net Borrowings - Interest * (1 - tax)

This is calculated below:

FCFE 0 1 2 3 4 5 6 7 8
Net Income 21125 21710 22295 22880 23465 24050 24635 27170
Add: Depreciation 23500 23500 23500 23500 23500 23500 23500 23500
Less: Change in WC 5000 0 0 0 0 0 0 -5000
Less: Capex 0 0 0 0 0 0 0 0
Add: Net Debt Availed 100000 -10000 -10000 -10000 -10000 -10000 -10000 -10000 -30000
Less: Interest * (1 - tax) 5850 5265 4680 4095 3510 2925 2340 1755
Asset Purchase -235000
FCFE -135000 23775 29945 31115 32285 33455 34625 35795 23915
NPV -16854
IRR 15.02%
Payback 4.53

Here, the Net Equity Outflow in Year 0 = Asset Purchase - Debt Availed = $ 235,000 - $ 100,000 = $ 135,000

Post that the FCFE is calculated for each year.

IRR

The IRR can be calculated using the IRR formula in excel i.e. = IRR (Range of cells)

The IRR comes out to 15.02%.

NPV

The NPV can be calculated using the excel NPV formula i.e. = NPV (rate, range of cells)

Here, the rate is given as 20 in the question.

Hence, = NPV (20%, Range of Cells)

The NPV is -$ 16,854

Note: If you notice, the  NPV is negative because the discount rate used is higher than the IRR (15%). Had the discount rate been lower, the NPV would be positive.

Payback Period

The payback of equity invested ($135,000) happens in approximately 5 years i.e. the overall investment is recovered in 5 years. This can be calculated as given below:

The sum of FCFE till Year 0 to 4 = $ -17,880 ...... (INR 17,880 still to be recovered)

The sum of FCFE till Year 0 to 5 = $ 15,575

FCFE in Year 5 = $ 33,455

This shows that the total cash flows turn positive between Year 4 and Year 5. Hence, taking a prorata assumption,

Out of the total FCFE in Year 5 ($33,455), the shortfall ($17,880) would be recovered in = 17,880 / 33455 years = 0.53 years

Hence, the absolute payback period would be = 4 + 0.53 years = 4.53 years